HUBBLE SPACE TELESCOPE DAILY REPORT #906 PERIOD COVERED: 0800 EST Daily Status Report as of 162/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed HSP/UV2 4943 (A Search for UV Variability in Nova 1992 - Cycle 2 High) The High Speed Photometer was used to attempt to make five ultraviolet observations of the supernova SN1992 Cygni (aka, GQ Muscae) to explore the physical conditions neat the white dwarf star in the system. All five images showed no sign of the target, with only a few reported pixels with one or two counts. 1.2 Completed S/C XSAFE (Safemode Recovery of the HST Science Instruments - FOC) The Safemode recovery of the Faint Object Camera was successful. 1.3 Completed WFC 4848 (GTO Low Latitude Parallel Program Cycle 3 High) The Wide Field Camera was used to make two low latitude sky survey observations taken in parallel with the prime science observations. Both observations were received and displayed in the OSS. They both were reported to look normal. 1.4 Completed FOC 4804 (UV Imaging of Nearby Galaxies with FOC/96) The Faint Object Camera was used to make one ultraviolet observations of nearby galaxy EAO-0308-5. The exposure lasted ten minutes and used filter F220W. The observation was started as planned with no problems, but completed during LOS. Final analysis will be completed after the playback of the science tape recorder data. 1.5 Completed GHRS/2 4067 (GHRS 1st Order Grating Spectral Lamp Test) The High Resolution spectrograph was used to perform a spectral lamp mini-function test for the first order grading. The observations occurred completely during an LOS, and the data remains on the science tape recorder. 1.6 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.7 Completed FOS/RD 4697 (FOS Polarmetric Calibration) Both the Red and Blue Sides of the Faint Object Spectrograph was used to obtain a series of measurements of the internal Pt-Ne lamp to locate the two oppositely polarized spectra that occur with the polarimeter. The observations occurred during LOS. The observations were monitored in the OSS, and no problems were reported with the telemetry. 1.8 Completed FOC/96 3913 (Ultraviolet Imaging of IRAS Starbursting Galaxies) The Faint Object Camera (f/48) was used to make ultraviolet image of starbursting galaxy IRAS12112+ using filters F140W and F220W. The observations were executed as scheduled, and no problems were reported. 1.9 Completed WFC 4381 (HST Medium Deep Survey - Cycle 3) The Wide Field Camera was used to make high latitude medium-deep sky survey observations taken in parallel with the prime science observations. The observations were monitored in the OSS, and the data quality was assessed. No problems were reported. 1.10 Completed FOS/RD 4062 (Stellar and Gas Dynamics in Normal Galaxies Augmentation: Cycle 2) The Faint Object Spectrograph (Red Side) was used to spectra of the inner one arcsecond of the core region of galaxy NGC-224 (aka, M-31, the Andromeda Galaxy). The observations were executed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 8 Successful : 8 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 7 Successful : 7 2.3 Significant HSTARs: None 2.4 Operations Note: Flight Software --------------- DF-224 Flight Software Version 9.0 software was successfully installed in the spacecraft as planned on June 9. The verification tests went well, and no major problems were encountered. This was the first time a complete new DF-224 version software was installed on the spacecraft since launch. SMS-161 was initiated at 161/0800Z and no major problems have been reported. Science observations are now continuing as normal. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Four Sets of PC 4714 (WF/PC Cycle 2 Calibration: Internal Delta Flats) 3.2 Complete Three Sets of WFC 4848 (GTO Low Latitude Parallel Program Cycle 3 High) 3.3 Complete Two Sets of FOC/96 4804 (UV Imaging of Nearby Galaxies with FOC/96) 3.4 Complete WFC 4929 (WP/PC Cycle 2 : Improved Dark Calibration) 3.5 Complete FOS/BL 4697 (FOS Polarimetric Calibration) 3.6 Complete WFC 4277 (Calibration of Nearby Galaxy Type IA Supernovae as Standard Candles) 3.7 Complete WFC/PC 4109 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor) 3.8 Complete FOS/RD 4062 (Stellar and Gas Dynamics in Normal Galaxies Augmentation: Cycle 2) 3.9 Complete Two Sets of FOC/96 3801 (Search for QSOs Suitable for Subsequent Observation fo He II) 3.10 Complete GHRS/2 4066 (GHRS Echelle Spectral Lamp Mini- Function Test) 3.11 Complete GHRS/2 4065 (Dark Noise Monitoring HRS) 3.12 Complete PC 4764 (Investigation of the Dynamics of Binary Asteroids) 3.13 Complete PC 4712 (WF/PC Cycle 2: Measles Monitor) 3.14 Complete FOS/BL 3441 (Post Asymptotic Giant Branch Evolution in the Magellanic Clouds) 3.15 Complete GHRS/2 3859 (Nighttime Measurements of Nitric Oxide - Cont. of 2544) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-